# Which Similarity-Sensitive Entropy (S-entropy)?

### v2

@@ Abstract
A canonical step in quantifying a system is to measure its entropy. Shannon entropy and other traditional entropy measures capture only the information encoded in the frequencies of a system’s elements. Recently, Leinster, Cobbold, and Reeve (LCR) introduced a method that also captures the far richer additional information encoded in the similarities and differences among elements, yielding similarity-sensitive entropy (S-entropy). More recently, the Vendi score (VS) was introduced as an alternative, raising the practical questions of how LCR and VS compare and which one to use. Here we address these questions conceptually, analytically, and experimentally, using 53 machine-learning datasets. We show that LCR and VS values can differ by orders of magnitude and can capture complementary information about a system, except in limiting cases. We demonstrate that both LCR and VS depend on how similarities are scaled and introduce the concept of “half distance” to parameterize this dependence. We prove that VS provides an upper bound on LCR for several values of the Rényi-Hill order parameter and present evidence that this bound holds for all values. We conclude that VS is preferable only when interpreting elements as linear combinations of a more fundamental set of “ur-elements” or when the system or dataset possesses a quantum-mechanical character. In the broader circumstance where one seeks simply to capture the rich information encoded by similarity as part of one’s entropy calculation, LCR is favored; nevertheless, for certain half-distances the two methods can complement each other.
@@ 1 Introduction
Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across science, engineering, commerce, and law [leinster2016maximizing, Shannon1951, Montemurro1994, tawilCompetitionCaliforniasMediCal2022]. Traditional entropy, as formulated by Shannon [shannon1948mathematical] and generalized by Rényi [renyiMeasuresEntropyInformation1961] (and later Tsallis [tsallis1988possible]), depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits or nats required to encode the shape of the distribution, with smaller values for more skew distributions and larger values for flatter ones (Fig. 1). These measures rely solely on frequencies and therefore ignore a rich additional source of information about the system: the pairwise similarities and differences among its elements (Fig. 1).
@@
Incorporating similarities into the calculation of entropy results in so-called similarity-sensitive entropy or S-entropy. This was first accomplished by Leinster and Cobbold [leinster2012measuring] and extended by Reeve and colleagues [reeve2014partition], in a framework we refer to here as LCR [leinsterEntropyDiversityAxiomatic2020]. By incorporating similarity, LCR can differentiate between systems that have identical frequency distributions but whose elements vary in how alike they are (Fig. 1). This capability is valuable in many domains [nguyen2023greylock], especially in machine learning (ML) [couch2024beyond, couch2025x] (for example, to achieve state-of-the-art performance more efficiently [chinnENRICHingMedicalImaging2023]), where datasets generally consist of all-unique elements, meaning any two datasets that are the same size will have the same traditional (i.e. similarity-insensitive) entropies because their frequency distributions are both flat. S-entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [arnaoutFutureBloodTesting2021, aroraRepertoirescaleMeasuresAntigen2022], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat and therefore traditional entropies less informative.
More recently, an alternative form of S-entropy has been described, consisting of the Vendi score [friedman2022vendi, pasarkar2025vendiscope] and its variants [pasarkar2023cousins]—hereafter collectively VS—separate from LCR. This raises the practical question of which form of S-entropy to choose for a given situation: LCR or VS. The current work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets used for ML, and comparing their values.
@@ 2 Related Work
Entropies fall into two classes: traditional/similarity‑insensitive and similarity‑sensitive. Although one often sees reference to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), entropies are actually families of measures; individual members are distinguished by how they weight frequencies and, for S-entropy, by their similarity matrix (see Table 1).
@@ 2.1 Traditional, similarity-insensitive entropy
Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the system [jost2006entropy]. Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ give greater weight to more frequent elements. $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form. Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [hill1973diversity]. For this reason, Hill’s formulation is often described as a unifying framework. $D_{q}$ are known as the Hill numbers or D numbers (for “diversity”) and are given by
@@
where $\mathbf{p}$ denotes the frequency distribution of unique elements. Other generalizations such as Tsallis entropy [tsallis1988possible] also exist. What these traditional entropies have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
2.2 Similarity-sensitive entropy (S-entropy)
@@ 2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
LCR [leinsterEntropyDiversityAxiomatic2020] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropy does not. A similarity matrix $Z$ is introduced, with entries $z_{ij}\in[0,1]$ that quantify the similarity between elements $i$ and $j$. The similarity of each element to itself is set to 1, making $Z$’s diagonal entries 1. The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity‑sensitive versions of the Rényi entropies $H_{\alpha}^{Z}$ and are given by
@@
Here $(Z\mathbf{p})_{i}=\sum_{j}z_{ij}p_{j}$ is the frequency‑weighted average similarity of element $i$ to all the elements (including itself). When this average is large, element $i$ is considered ordinary, making Eq. 2 interpretable as the average “ordinariness” [leinster2012measuring] across all elements. $D_{q}^{Z}{(\mathbf{p})}$ appears widely in the recent literature, where it is variously known as:
  • phylogenetic diversity [chao2010phylogenetic] in the special case where $z_{ij}$ forms an ultrametric, typically derived from a phylogenetic tree;
  • functional diversity [chaoAttributediversityApproachFunctional2019] when the similarity pertains to elements’ function, for example the binding similarity between pairs of antibodies or TCRs [aroraRepertoirescaleMeasuresAntigen2022];
  • attribute diversity [chaoAttributediversityApproachFunctional2019] when interpreting the system as a set of attribute contributions instead of frequencies; and
  • similarity-sensitive or similarity-aware diversity more generally [nguyen2023greylock].
LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become uninformative (Fig. 1). Representative applications in the life sciences include high‑throughput immunology (immunomes) [aroraRepertoirescaleMeasuresAntigen2022], microbiome research (metagenomics) [nguyen2023greylock], and medical imaging [couch2024beyond] Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g. images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [couch2024beyond].
2.2.2 The Vendi score and its “cousins” (VS)
@@
VS is to this similarity matrix what the von Neumann entropy is to the quantum density matrix. Replacing the Shannon term in Eq. 4 with a Rényi entropy of order $q$ yields the so‑called “cousins” [pasarkar2023cousins] of the Vendi score:
@@ 3.1 Datasets and definitions
  • Image datasets. 22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [yangMedMNISTClassificationDecathlon2021a]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [visipediaInat_comp2017READMEmd]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [mei2022radimagenet], the NIH ChestXRay dataset [wangChestXRay8HospitalScaleChest2017]; COCO (Common Objects in Context) [linMicrosoftCOCOCommon2015]; and MIDRC COVIDx CXR-4 [wuCOVIDxCXR4Expanded2023]. For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset were downscaled to $250\times 250$ pixels, and images in the COCO datast were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
  • Tabular datasets. After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [lucieImprovedPython2024]. For each dataset, identifier columns and any features containing non‑numeric strings or time‑series data were removed.
@@ 3.2 Similarity matrices
where $k$ is a parameter that was varied during experiments to test for effects on results. This similarity measure is provably positive‑definite (PSD), guaranteeing valid similarity matrices for VS (see Section 6); LCR has no such restriction [leinsterEntropyDiversityAxiomatic2020]. Imaging datasets were first embedded into a two-dimensional space using UMAP [mcinnesUMAPUniformManifold2020] and the default $k$ taken to be $2^{-1/2}$, making the norm in Eq. 7 equivalent to the root-mean-square distance (RMSD). The default $k$ for tabular datasets was $1$.
3.3 Entropy calculations
The greylock Python package [nguyen2023greylock] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section 2.1). The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with the default VS (see Section 2.2.2). VS was calculated using the vendi‑score package [vendi-score]. This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR. Similarity matrices were calculated on a GPU-enabled workstation equipped with multi‑core CPUs. LCR and VS computations were performed on a 2024 M4 Apple Mac mini with 64GB unified memory.
@@ 3.4 Clustering
HDBSCAN (Hierarchical Density‑Based Spatial Clustering of Applications with Noise) [mcinnesHdbscanHierarchicalDensity2017] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset. HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well‑suited for heterogeneous data.
@@ 3.5 Numerical experiments
Following a previous approach [Couch:2023pav], we searched for counterexamples to Conjecture 1 (Section 4.3) by numerically minimizing $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for uniform $\mathbf{p}$ using scipy.optimize. A negative value indicates a counterexample. To meet the assumptions of this conjecture, $Z$ was constructed as a gram matrix (which are always PSD), finding an $r\times n$ matrix $w$, for some integer $r$, such that $Z=w^{\intercal}w$. To ensure that the diagonal elements of $Z$ are equal to $1$, the rows of $w$ were $r$-dimensional unit vectors; to ensure the entries of $Z$ are non-negative, we took the entries of $w$ to be non-negative. We assumed a uniform distribution of unique elements. Thus, we parameterized $Z$ by a $r\times n$ matrix $x$, from which we defined
@@ 4.1 LCR and VS of major imaging and tabular ML datasets
We began by comparing LCR to VS, both at $q=1$, on a simple, well‑known imaging dataset: MNIST handwritten digits; the results below are for a random 10,000-image subset, taken for computational convenience. Both quantities were expressed in their effective‑number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state‑of‑the‑art clustering algorithm HDBSCAN.
@@
Fig. 2a displays a UMAP embedding of the data, coloring each digit by its label. HDBSCAN detects six clusters, grouping together digits that appear visually similar (Fig. 2b). LCR yields the effective number of unique images after accounting for pairwise image similarity; it allows visually similar images to each contribute less to the entropy than if they were completely distinct (Fig. 2c). Consequently, points that lie close together count less toward the total, whereas points separated by some distance contribute more independently, even for points that happen to be in the same cluster. Taken together, these considerations suggest that LCR will approximate the number of clusters—potentially exceeding it slightly because the clusters are somewhat loose. Indeed, the similarities among the images in this 10,000-image subset result in it having an effective number of just 12.5 images according to LCR: roughly twice the number of clusters and 25% greater than the number of classes (10). Mathematically, this indicates that this MNIST subset possesses the same entropy as a hypothetical dataset containing 12-13 completely dissimilar elements (i.e. with zero pairwise similarity). This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the S-entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease the entropy relative to a scenario in which all 10 digit classes were entirely distinct.
In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements. However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen‑ or “ur”‑images, each actual image in the dataset being a linear combination of these. Fig. 3a displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy (Fig. 3b). Some of these eigen‑images resemble recognizable digits, whereas others resemble superpositions of multiple digits. VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen‑images at frequencies given by the eigenvalues. The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small.
@@
Fig. 4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets.

### v3

@@ Abstract
Shannon entropy is not the only entropy that is relevant to machine‑learning datasets, nor possibly even the most important one. Traditional entropies such as Shannon entropy capture information represented by elements’ frequencies but not the richer information encoded by their similarities and differences. Capturing the latter requires similarity‑sensitive entropy (S‑entropy). S‑entropy can be measured using either the recently developed Leinster–Cobbold–Reeve framework (LCR) or the newer Vendi score (VS). This raises the practical question of which one to use: LCR or VS. Here we address this question conceptually, analytically, and experimentally, using 53 large and well-known imaging and tabular datasets. We find that LCR and VS values can differ by orders of magnitude and are complementary, except in limiting cases. We show that both LCR and VS results depend on how similarities are scaled, and introduce the notion of “half‑distance” to parameterize this dependence. We prove that VS provides an upper bound on LCR for several values of the Rényi–Hill order parameter and present evidence that this bound holds for all values. We conclude that VS is preferable only when a dataset’s elements can be usefully interpreted as linear combinations of a more fundamental set of “ur‑elements” or when the system that the dataset describes has a quantum‑mechanical character. In the broader case where one simply wishes to capture the rich information encoded by elements’ similarities and differences as well as their frequencies, LCR is favored; nevertheless, for certain half‑distances the two methods can complement each other.
@@ 1 Introduction
Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across not only science and engineering but in commerce and law [15, 29, 22, 31]. Traditional entropy, as formulated by Shannon [30] and generalized by Rényi [28] (and later Tsallis [32]), depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits (or nats) required to encode the shape of the distribution, with skew distributions requiring fewer bits and flatter distributions requiring more (Fig. 1). These entropies rely solely on frequencies and therefore ignore a rich additional source of information about the system: the pairwise similarities and differences among its elements (Fig. 1).
@@
Incorporating similarity into the calculation results in so-called similarity-sensitive entropy or S-entropy. This was first accomplished by Leinster and Cobbold [14] and extended by Reeve and colleagues [27], in a framework we refer to as LCR [17]. By incorporating similarity, LCR can differentiate between systems that have identical frequency distributions but whose elements vary in how alike they are (Fig. 1). This capability is valuable in many domains [23], especially in machine learning (ML) [6, 7] (for example, to achieve state-of-the-art performance more efficiently [5]), where datasets generally consist of all-unique elements, meaning any two datasets that are the same size will have the same traditional (i.e. similarity-insensitive) entropies because their frequency distributions are both flat, even though they may differ dramatically in composition. S-entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [1, 2], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat, such that traditional entropies are less informative.
More recently, an alternative form of S-entropy to LCR has been described, consisting of the Vendi score [9, 26] and its variants or “cousins” [25]—hereafter collectively VS. This alternative raises the practical question of which form of S-entropy to choose for a given system or dataset: LCR or VS. The present work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets commonly used for ML research, and comparing their values.
@@ 2 Related Work
Entropies fall into two classes: (i) traditional, i.e. similarity‑insensitive, and (ii) similarity‑sensitive, a.k.a. S-entropy. Although the literature often refers to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), both classes are actually families of measures; individual members are distinguished by how they weight frequencies and, for S-entropy, by their similarity matrix as well (see Table 1).
@@ 2.1 Traditional, similarity-insensitive entropy
Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the dataset or system [13]. Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ give greater weight to more frequent elements. $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form. Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [12]. For this reason, Hill’s formulation is often described as a unifying framework. $D_{q}$ are known as the Hill numbers or D numbers (for “diversity”) and are given by
@@
$\mathbf{p}$ denotes the frequency distribution of unique elements. Other generalizations such as Tsallis entropy [32] also exist. What these traditional entropies have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
2.2 S-entropy
@@ 2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
LCR [17] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropies do not. A similarity matrix $Z$ is introduced, with entries $z_{ij}\in[0,1]$ that quantify the similarity between elements $i$ and $j$. The similarity of each element to itself is set to 1, making $Z$’s diagonal entries 1. The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity‑sensitive versions of the Rényi entropies $H_{\alpha}^{Z}$ and are given by
@@
Here $(Z\mathbf{p})_{i}=\sum_{j}z_{ij}p_{j}$ is the frequency‑weighted average similarity of element $i$ to all the elements (including itself). When this average is large, element $i$ is considered ordinary, making Eq. 2 interpretable as the average “ordinariness” [14] across all elements. $D_{q}^{Z}{(\mathbf{p})}$ appears widely in the recent literature, where it is variously known as:
  • phylogenetic diversity [3] in the special case where $z_{ij}$ forms an ultrametric, typically derived from a phylogenetic tree;
  • functional diversity [4] when the similarity pertains to elements’ function, for example the binding similarity between pairs of antibodies or TCRs [2];
  • attribute diversity [4] when interpreting the system as a set of attribute contributions instead of frequencies; and
  • similarity-sensitive or similarity-aware diversity more generally [23], for which “S-entropy” is a convenient shorthand.
LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become less informative (Fig. 1). Representative applications in the life sciences include high‑throughput immunology (immunomes) [2], microbiome research (metagenomics) [23], and medical imaging [6] Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g. images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [6].
2.2.2 The Vendi score and its variants (VS)
@@
VS is to this similarity matrix what the von Neumann entropy is to the quantum density matrix. Replacing the Shannon term in Eq. 4 with a Rényi entropy of order $q$ yields the so‑called “cousins” [25] of the Vendi score:
@@ 3.1 Datasets and definitions
  • Image datasets. 22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [36]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [33]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [21], the NIH ChestXRay dataset [34]; COCO (Common Objects in Context) [18]; and MIDRC COVIDx CXR-4 [35]. For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset were downscaled to $250\times 250$ pixels, and images in the COCO datast were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
  • Tabular datasets. After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [11]. For each dataset, identifier columns and any features containing non‑numeric strings or time‑series data were removed.
@@ 3.2 Similarity matrices
where $k$ is a parameter that was varied during experiments to test for effects on results. This similarity measure is provably positive‑definite (PSD), guaranteeing valid similarity matrices for VS (see Section 6); LCR has no such restriction [17]. Imaging datasets were first embedded into a two-dimensional space using UMAP [20] and the default $k$ taken to be $2^{-1/2}$, making the norm in Eq. 7 equivalent to the root-mean-square distance (RMSD). The default $k$ for tabular datasets was $1$.
3.3 Entropy and S-entropy calculations
The sentropy Python package [23] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section 2.1). The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with the default VS (see Section 2.2.2). VS was calculated using the vendi‑score package [10]. This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR. Similarity matrices were calculated on a GPU-enabled workstation equipped with multi‑core CPUs. LCR and VS computations were performed on a 2024 M4 Apple Mac mini with 64GB unified memory.
@@ 3.4 Clustering
HDBSCAN (Hierarchical Density‑Based Spatial Clustering of Applications with Noise) [19] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset. HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well‑suited for heterogeneous data.
@@ 3.5 Numerical experiments
Following a previous approach [8], we searched for counterexamples to Conjecture 1 (Section 4.3) by numerically minimizing $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for uniform $\mathbf{p}$ using scipy.optimize. A negative value indicates a counterexample. To meet the assumptions of this conjecture, $Z$ was constructed as a gram matrix (which are always PSD), finding an $r\times n$ matrix $w$, for some integer $r$, such that $Z=w^{\intercal}w$. To ensure that the diagonal elements of $Z$ are equal to $1$, the rows of $w$ were $r$-dimensional unit vectors; to ensure the entries of $Z$ are non-negative, we took the entries of $w$ to be non-negative. We assumed a uniform distribution of unique elements. Thus, we parameterized $Z$ by a $r\times n$ matrix $x$, from which we defined
@@ 4.1 LCR and VS of major imaging and tabular ML datasets
We began by comparing LCR to VS, both at $q=1$, on a simple, well‑known imaging dataset: MNIST handwritten digits. The results below are for a random 10,000-image subset, taken for computational convenience. Both quantities were expressed in their effective‑number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state‑of‑the‑art clustering algorithm HDBSCAN.
@@
Fig. 2a displays a UMAP embedding of the data, coloring each digit by its label. HDBSCAN detects six clusters, successfully grouping together digits that appear visually similar (Fig. 2b). LCR yields the effective number of unique images after accounting for pairwise image similarity; it allows visually similar images to each contribute less to the entropy than if they were completely distinct from each other (Fig. 2c). Consequently, points that lie close together count less toward the total, whereas points separated by some distance contribute more independently, even for points that happen to be in the same cluster. Taken together, these considerations suggest that LCR should approximate the number of clusters—potentially exceeding it slightly because the clusters are somewhat loose. Indeed, the similarities among the images in this 10,000-image subset result in it having an effective number of just 12.5 images according to LCR: roughly twice the number of clusters and 25% greater than the number of classes (10). Mathematically, this indicates that this MNIST subset possesses the same entropy as a hypothetical dataset containing 12-13 completely dissimilar elements (i.e. with zero pairwise similarity). This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the S-entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease it relative to a scenario in which all 10 digit classes were entirely distinct.
In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements. However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen‑ or “ur”‑images, each actual image in the dataset being a linear combination of these. Fig. 3a displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy (Fig. 3b, gray). Some of these eigen‑images resemble recognizable digits, whereas others resemble superpositions of multiple digits. VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen‑images at frequencies given by the eigenvalues. The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small, as confirmed by Fig. 3b.
@@
Fig. 4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets, demonstrating large differences for imaging datasets and smaller (and sometimes no) differences for tabular datasets.